OpenStudy (anonymous):

x^2+4x+4x+16=8 enter the quadratic equation in standard form.

OpenStudy (anonymous):

x^2+8x+8=0

OpenStudy (anonymous):

The standard form of any quadratic equation is \[ax^2+bx+c=0\]

OpenStudy (anonymous):

Where a,b,c are constants

OpenStudy (anonymous):

hey the left side can be factorised as (x+4)^2=8 x+4=\[\[\pm\]x+4=2\sqrt{2}\]

OpenStudy (anonymous):

ofcourse a cannot be equal to 0

OpenStudy (anonymous):

sry typing error \[x+4=\pm \sqrt{8}\]

OpenStudy (anonymous):

@vamgadu , I think @Astemmene is looking for standard form of quadratic equation.

OpenStudy (anonymous):

oh i dint look and you caught me!!:P

OpenStudy (anonymous):

np :)